# Axial Bianchi I meets drifting extragalactic sources

### v1

@@ Abstract
We recompute the drift of comoving sources in axial Bianchi I universes and correct the same calculation by Quercellini et al. from 2009. The correction is pertinent with respect to the Gaia catalogue of measured quasar drifts, that beg to be fitted.
@@ 1 Introduction
As early as 1933, Georges Lemaître [1], following a recommandation by Albert Einstein, considers (tri-axial) Bianchi I universes without using this name. They generalize flat Robertson-Walker universes by allowing three different scale factors in three orthogonal directions:
@@
The proper time element of the axial Bianchi I universe is obtained by choosing two equal scale factors, e.g. $b=a$. Then there is a fourth Killing vector, $x\partial_{y}-y\partial_{x}$, generating rotations around the $z$-axis. If all three scale factors are equal, we are back at flat Robertson-Walker universes with the maximally symmetric Euclidean 3-space and three translations, three rotations.
The purpose of this work is an ab initio calculation of the drift of a comoving source as seen by a not necessarily comoving observer, whose peculiar velocity however is much smaller than the speed of light. We carefully separate the three parts of the calculation:
kinematics:
position of the source on the celestial sphere as a function of the observer’s cosmic time and the redshift of the source,
dynamics:
Einstein’s equation with cosmological constant and comoving dust,
approximation of small anisotropy:
$c(t)=a(t)[1+{\textstyle\frac{3}{2}}\eta(t)],\ \eta(t)\ll 1$.
Note that only axial Bianchi I and IX universes fulfill the three conditions: (i) next-to-maximal symmetry, (ii) compatibility with Einstein’s equation sourced by comoving dust and (iii) admitting maximal symmetry in the limit $\eta\rightarrow 0$.
In 2014 Darling [6] used tri-axial Bianchi I universes to fit the drift of 429 extra-galactic radio sources measured by Titov & Lambert [7] using Very Long Baseline Interferometry.
@@ 2 The path of a light ray
We follow notations and use results of [8]. Consider a geodesic $x^{\mu}(p)$ and denote the derivation with respect to the affine parameter $p$ by ˙:$\dot{}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}p$. The four Killing vectors $\partial x,\,\partial y,\,\partial z$ and $x\partial y-y\partial x$ yield four conserved quantities by Emmy Noether’s theorem:
$A\vcentcolon=a(t(p))^{2}\dot{x},\ B\vcentcolon=a^{2}\dot{y},\ C\vcentcolon=c^{2}\dot{z}\quad\text{and}\quad Ay-Bx=0.$ (3)
@@
Eliminating the affine parameter $p$ in favour of the cosmic time $t$ and using :${}^{\prime}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}t$, the light-like geodesic must satisfy the three first order equations:
@@
In our conventions the speed of light is unity; proper time $\tau$, the coordinate time $t$ and the affine parameter $p$ have units of seconds, the comoving coordinates $x,\,y,\,z$ are dimensionless and the scale factors $a$ and $c$ and the Noether constants $A,\,B$ and $C$ carry seconds. Then $W$ is dimensionless. Equation (5) remains unchanged under a simultaneous rescaling of the three Noether constants. Therefore we may choose $N\vcentcolon=\sqrt{A^{2}+B^{2}+C^{2}}$ to be any positive constant. Through a rescaling of the coordinates $x,\,y,\,z$ we may achieve $a_{0}=c_{0},$ with $a_{0}\vcentcolon=a(t_{0}),\dots$
@@ 3 A second, infinitesimally close light ray
    If we pretend to know the metric and it has nice properties, then the redshift allows us to compute the time of flight of the photon.
  • $T_{D}$ is a few years and we observe a change in the direction of the arriving photons. Of course, a change in direction can only be measured with respect to another direction that supposedly does not change. Axial Bianchi I universes have indeed the luxury of such fixed directions: the $z$-axis and the $xy$-plane.
@@
$\vec{V}=\vcentcolon a_{0}\,\vec{v}\,,\quad\quad V\ll 1.$ (9)
@@
We suppose that $T$ is small with respect to the ‘time of flight’, $T\ll t_{0}-t_{e}$; so small that we may keep only linear terms in $T$ and $T_{D}$. Also the peculiar velocity is small enough so that we may keep only linear terms in $V$. In particular, we neglect the dilation of the observer’s proper time with respect to cosmic time $t$.
@@
Our task is to slightly change the initial direction of the second geodesic $\tilde{\vec{x}}(t)$ coded in the Noether constants $\tilde{A},\,\tilde{B}$ and $\tilde{C}$, in such a manner that $\tilde{\vec{x}}_{\tilde{0}}=\vec{x}_{0}+T_{D}\,\vec{v}$. To this end we set:
$\tilde{A}=A+\alpha N,\quad\tilde{B}=B+\beta N,\quad\tilde{C}=C+\gamma N,\quad\alpha,\,\beta,\,\gamma\sim T/(t_{0}-t_{e}).$ (12)
We may choose $\tilde{N}=N$ implying
@@
Then we have to first order in $T/(t_{0}-t_{e})$:
$\tilde{x}_{\tilde{0}}\!\!\!$ $\sim$ $\!\!\!\int_{t_{e}+T}^{t_{0}+T_{D}}\left(\frac{\tilde{A}\tilde{W}}{a^{2}}\,+\,\frac{NW}{a^{2}}\,\alpha\right)$ (20)
@@
$\sim$ $\!\!\!x_{0}+A\,F\,T_{D}+\left(\frac{AC}{A^{2}+B^{2}}\,\gamma+\alpha\right)I-\,\frac{AC}{A^{2}+B^{2}}\,\gamma J.$ (22)
@@
$\underline{z}_{0}+1\vcentcolon=\,\frac{T_{D}}{T}\,\sim\,\frac{1}{W_{e}}\,\frac{1}{N/a_{0}-(Av_{x}+Cv_{z})}\,.$ (28)
@@ 4 Arriving angles
Our first task is to compute the angle $\theta$ between the $z$-axis and the velocity $\vec{x}^{\prime}_{0}$ of the first photon. To this end we use the negative of the spatial part of the metric (2). Note that the velocity vector $\vec{x}^{\prime}$ is a unit vector with respect to this 3-metric, because we have set the speed of light to one. If we denote by
@@
Our second task is to compute the same angle $\tilde{\theta}$ for the second photon at its arrival time $t_{\tilde{0}}=t_{0}+T_{D}$:
@@
$\,\frac{\tilde{C}\tilde{W}}{c}\,\sim\,\frac{CW}{c}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW}{c}\right\}^{2}\right)\right],$ (36)
@@
$\,\frac{\tilde{C}\tilde{W}_{\tilde{0}}}{c_{\tilde{0}}}\,\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW_{0}}{c_{0}}\right\}^{2}\gamma\right)\right]\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N}{C}\,\gamma\right].$ (37)
@@
Note that if $A=B=0$ the first photon moves along the $z$-axis, $\sin\theta=0$. Then the projection of the photon’s velocity onto the $xy$-plane vanishes and the angle $\varphi$ is not defined.
@@ 5 Linearizing the scale factors
We known that the anisotropy of the universe cannot be large and we will concentrate on axial Bianchi I universes close to an underlying maximally symmetric one with only one scale factor $a_{F}(t)$, $\cdot_{F}$ for Friedman. It is convenient to parameterize the small deviation $\eta(t)$ in the following way:
$a=\vcentcolon a_{F}\,[1-{\textstyle\frac{1}{2}}\eta],\quad c=\vcentcolon a_{F}\,[1+\eta],\quad\quad\quad\eta_{0}=0,\quad H\sim H_{F}-{\textstyle\frac{1}{2}}\eta^{\prime},\quad H_{c}\sim H_{F}+\eta^{\prime},$ (50)
where we indicate the linearization in $\eta$ and $V$ (neglecting cross-terms $\eta V$) also by $\sim$. Then we have
@@
$\alpha$ $\sim$ $\frac{T_{D}}{a_{F0}\chi_{e0}}\,(3\,\cos\varphi\,\sin\theta\,\cos^{2}\theta\,\eta_{e}+[1-\cos^{2}\varphi\,\sin^{2}\theta]\,V_{x}$ (55)
@@
$\gamma$ $\sim$ $-\sin\theta\,\frac{T_{D}}{a_{F0}\chi_{e0}}\left(3\,\sin\theta\,\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right),$ (56)
@@
$\chi_{e0}\vcentcolon=\int_{t_{e}}^{t_{0}}\,\frac{1}{a_{F}}\,\quad\text{and}\quad\vec{V}=a_{F0}\vec{v}.$ (57)
Consider first the drift with respect to $\varphi$, equation (49), and linearize it in $\eta$ and $V$, but not in $T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}$:
$\cos\tilde{\varphi}\sim\hskip 409.71968pt$ (58)
$\frac{\cos\varphi+[3\cos\varphi\,\cos^{2}\!\theta\,\eta_{e}+(1-\cos^{2}\!\varphi\,\sin^{2}\!\theta)\!\!\left\{V_{x}/\sin\theta\right\}-\cos\varphi\,\cos\theta\,V_{z}]\,T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}}{\sqrt{1+2\cos\theta[3\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\left\{V_{x}/\sin\theta\right\}-V_{z}]\,T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}}}\,.$
@@
$\cos\tilde{\varphi}$ $\sim$ $\cos\varphi\left[1+\,\frac{\sin^{2}\varphi}{\cos\varphi}\,\frac{V_{x}}{\sin\theta}\,\frac{T_{D}}{a_{F0}\chi_{e0}}\right].$ (60)
@@
$\delta\varphi$ $\sim$ $\frac{-1}{\sin\varphi}\,\,\delta\cos\varphi\sim-\,{\sin\varphi}\,\frac{V_{x}}{\sin\theta}\,\frac{T_{D}}{a_{F0}\chi_{e0}}\,.$ (62)
@@
$\cos\tilde{\theta}$ $\sim$ $\cos\theta\left[1-\sin^{2}\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\,\right)\frac{T_{D}}{a_{F0}\chi_{e0}}\right],$ (63)
$\delta\cos\theta$ $\sim$ $-\sin^{2}\theta\,\cos\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\right)\frac{T_{D}}{a_{F0}\chi_{e0}},$ (64)
$\delta\theta$ $\sim$ $\left(3\sin\theta\,\cos\theta\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right)\,\frac{T_{D}}{a_{F0}\chi_{e0}}\,.$ (65)
@@
Our result disagrees with the result by Quercellini et al. [10][11] without peculiar velocity. They find – as the redshift ranges from zero to one – that the absolute value of the ratio is at most 7 %. They also find the opposite sign in front of the first term. The sign problem was noticed before by Marcori et al. [12], who also include peculiar velocities.

### v2

@@ Abstract
We recompute ab initio the drift of comoving sources in axial Bianchi I universes. Our result resolves the long-standing disagreement between the computations by Quercellini et al. from 2009 and by Marcori et al. from 2018 in favour of the latter. These computations have a potential impact on Gaia data analysis and cosmological parameter estimation.
@@ 1 Introduction
As early as 1933, Georges Lemaître [1], following a recommendation by Albert Einstein, considers (tri-axial) Bianchi I universes without using this name. They generalize flat Robertson-Walker universes by allowing three different scale factors in three orthogonal directions:
@@
The proper time element of the axial Bianchi I universe is obtained by choosing two equal scale factors, e.g. $b=a$. Then there is a fourth Killing vector, $x\partial_{y}-y\partial_{x}$, generating rotations around the $z$-axis. Note that if all three scale factors are different, the Bianchi I universe has three privileged axes, the $x$-, $y$- and $z$-axis. The axial Bianchi I universe has one privileged axis. If all three scale factors are equal, we are back at flat Robertson-Walker universes with the maximally symmetric Euclidean 3-space, three translations, three rotations and no privileged axis.
1.1 Drift of extra-galactic sources in Bianchi I universes
  • In 2009 Quercellini et al. [6] compute the drift of comoving sources as seen by a comoving observer in tri-axial Bianchi I universes. It is given by their equations (6) and (7). See also Quercellini et al. [7], equations (33) and (34).
  • In 2013 Titov & Lambert [8] measure the drift of 429 extra-galactic radio sources using Very Long Baseline Interferometry.
  • In 2014 Darling [9] applies the drift equations (6) and (7) by Quercellini et al. to fit the drift of the radio sources observed by Titov & Lambert. He finds no significant signal.
  • In 2018 Marcori et al. [10] recompute the drift including peculiar velocities of source and observer. Although they do not spell out the drift equations their calculation is in disagreement with equations (6) and (7) by Quercellini et al..
@@ 1.1 Drift of extra-galactic sources in Bianchi I universes
1.2 Aim and motivation of choices
The purpose of this work is resolving the mentioned disagreement and presenting correct drift equations ready to be confronted with the new high-quality Gaia data. Our calculation is ab initio and low tech and concerns the drift of a comoving source in axial Bianchi I universes filled with comoving dust and admitting small anisotropy. We also include an observer’s small peculiar velocity.
Let us comment our choices:
ab initio:
Our only prerequisites are the geodesic equation, Einstein’s equation and the definition of cosmic time $t$.
low tech:
Our aim is that an experimental colleague can verify all calculations.
As Quercellini et al. we work in coordinates and use Cartesian coordinates $(x,y,z)$ and their spherical coordinates $(r,\varphi,\theta)$.
All approximations are first order Taylor expansions in four small dimensionless quantities:
  • Light of atomic period $T$ is emitted by an extragalactic source and observed with Doppler-shifted period $T_{D}$, spectroscopic redshift and incoming direction. We suppose that its time of flight $t_{of}$ is superior to that of light from Andromeda $t_{of}>10^{14}\,$s. Admitting the detection limit of the James Webb Space Telescope, $T_{D}<10^{13}\,$s, we define the small dimensionless parameter $\epsilon_{s}\vcentcolon=T_{D}/t_{of}<10^{-27}$, $\cdot_{s}$ for spectroscopic, and use it to compute the redshift by means of a first order Taylor expansion in $\epsilon_{s}$.
  • The incoming direction is recorded over a tracking time of say 10 years, which we also denote by $T_{D}$. We define $\epsilon_{d}\vcentcolon=T_{D}/t_{of}<3\cdot 10^{-6}$, $\cdot_{s}$ for drift.
  • We allow the observer to have a small peculiar velocity $V$ divided by the speed of light (that we set to 1). We will restrict the fit to $V<10\,\%.$
  • We allow for a small time-dependent anisotropy, $\eta(t)$ with $|\eta(t)|<10\,\%$ between emission and reception today, $t\in[t_{e},t_{0}]$.
These four small quantities are considered independent variables, not necessarily of the same order of magnitude. The drift equations e.g.  (66) are of the form $\delta\theta\sim(a\,\eta+b\,V)\,\epsilon_{d}$. We denote by the symbol $\sim$ an equality neglecting terms in $\epsilon_{d}^{2}$, $\eta^{2}$, $V^{2}$ or $\eta\,V$ and higher orders, such that each neglected term is less than 1 %. We suppose that all functions encountered are analytic and acknowledge that the domains of convergence must be controlled during the fit by reducing the domain of input parameters as necessary. In the example we must control that the coefficients $a$ and $b$ remain of order 1.
comoving source:
Peculiar velocities of the sources are not included because they would add three more parameters for each single source. But we may hope that in a large enough sample the peculiar velocities of the sources average out and become less and less important with increasing redshift.
observer’s small peculiar velocity:
A small peculiar velocity of the observer is considered because it only adds three more parameters, which might be relevant for sources with redshifts below 0.5.
Axial
Bianchi universes have four Killing vectors while generic Bianchi universes have only three Killing vectors. In this sense the axial Bianchi universes are closer to the ‘Robertson-Walker’ universes that obey the cosmological principle (six Killing vectors) than the generic Bianchi universes. Now the four Killing vectors of axial Bianchi I universes imply geodesics with four conserved quantities, which in turn produce equations for redshift and drift that are much simpler and involve less arbitrary parameters in their fit compared to tri-axial Bianchi I universes.
Nevertheless axial Bianchi I universes are sufficient to trace the errors in Quercellini et al.’s calculation.
small anisotropy:
As the cosmological principle with its six space-like Killing vectors is quite successful in fitting today’s data it is a wise choice to remain close to Robertson-Walker universes. For axial Bianchi I universes this means small anisotropy $\eta(t)\vcentcolon={\textstyle\frac{2}{3}}[c(t)/a(t)-1]\rightarrow 0$. We will be more precise about this parametrization and motivate it at the beginning of section 5. Note that there are only two other axial universes admitting this limit: axial Bianchi V and IX. Axial Bianchi V however is incompatible with Einstein’s equation sourced by comoving dust [11][12].
During our calculation we carefully separate the three parts: (i) kinematics: position of the source on the celestial sphere as a function of the observer’s cosmic time and the redshift of the source, (ii) dynamics: Einstein’s equation with cosmological constant and comoving dust, (iii) approximation of small anisotropy: $|\eta(t)|\ll 1$.
On the other hand Quercellini et al. mix kinematics and dynamics in their calculation.
1.3 Summary of results
Quercellini et al. use polar angles $\varphi$ and $\theta$ to compute the drift of a comoving source as seen today $t_{0}$ by a comoving observer, $V=0$, in tri-axial Bianchi I universes. Neglecting the curvature of the photons’ trajectories they find equations (6) and (7) (in [6]) for the $\varphi$- and $\theta$- components of the drift. Note that these equations are independent of the photon’s emission time and of its redshift. It is claimed that the approximation is good to 7 % for redshifts up to 1, if the Bianchi I universe satisfies Einstein’s equations with positive cosmological constant $\Lambda$ and comoving dust today $\rho_{0}$.
In the axial case, $b=a$, the $\varphi$-drift vanishes and one is left with equation (7), which becomes independent of $\varphi$.
For small anisotropy $\eta(t)$ and to linear order in $\eta$, their equation (7) reduces further to
$\delta\theta\ \sim\ 3\sin\theta\cos\theta\left(-{\textstyle\frac{1}{2}}\,\eta^{\prime}_{0}\quad\quad\quad\quad\quad\quad\right)\,T_{D}\,,$
$T_{D}$ being the duration of observation. Our calculation yields (equation (66) below with $V=0$),
$\delta\theta\ \sim\ 3\sin\theta\cos\theta\left(+{\textstyle\frac{1}{2}}\,\eta^{\prime}_{0}+\eta_{e}/a_{F0}/\chi_{e0}\right)\,T_{D}\,,$
with the emission time $t_{e}$ and
$\eta_{e}\vcentcolon=\eta(t_{e}),\quad\chi_{e0}\vcentcolon=\int_{t_{e}}^{t_{0}}1/a_{F}\vcentcolon=\int_{t_{e}}^{t_{0}}1/a_{F}(t)\,\hbox{\rm{d}}t.$
The emission time $t_{e}$ is computed from the direction-dependent redshift $\underline{z}_{0}$ using our equation (53) with $V=0$,
$\underline{z}_{0}+1\ \sim\ \frac{a_{F0}}{a_{F}(t_{e})}[1+{\textstyle\frac{1}{2}}(1-3\cos^{2}\theta)\eta_{e}].$
We explain in equation (69) that for $\underline{z}_{0}=1$ the redshift dependent piece, which is neglected by Quercellini et al., is in absolute value more than four times larger than the redshift independent piece, rather than smaller by a factor of 7 % as claimed by the authors. Notice that the Gaia catalogue contains objects with redshift up to 4.7.
As shown we also find the opposite sign in front of the redshift independent piece.
Our results and those of Marcori et al. [10] agree in their domain of overlap.
@@ 2 The path of a light ray
We follow the notations of [13]. Consider a geodesic $x^{\mu}(q)$ and denote the derivation with respect to the affine parameter $q$ by ˙:$\dot{}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}q$. The four Killing vectors $\partial_{x},\,\partial_{y},\,\partial_{z}$ and $x\partial_{y}-y\partial_{x}$ yield four conserved quantities by Emmy Noether’s theorem:
$A\vcentcolon=a(t(q))^{2}\dot{x},\ B\vcentcolon=a^{2}\dot{y},\ C\vcentcolon=c^{2}\dot{z}\quad\text{and}\quad Ay-Bx=0.$ (3)
The physical significance of the Noether constants $A$, $B$ and $C$ in terms of conserved quantities is summarized in appendix A.
@@
Eliminating the affine parameter $q$ in favour of the cosmic time $t$ and using :${}^{\prime}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}t$, the light-like geodesic must satisfy the three first order equations:
@@
In our conventions the speed of light is unity; proper time $\tau$, the coordinate time $t$ and the affine parameter $q$ have units of seconds, the comoving coordinates $x,\,y,\,z$ are dimensionless and the scale factors $a$ and $c$ and the Noether constants $A,\,B$ and $C$ carry seconds. Then $W$ is dimensionless. Equation (5) remains unchanged under a simultaneous rescaling of the three Noether constants. Therefore we may choose $N\vcentcolon=\sqrt{A^{2}+B^{2}+C^{2}}$ to be any positive constant. Through a rescaling of the coordinates $x,\,y,\,z$ we may achieve $a_{0}=c_{0},$ with $a_{0}\vcentcolon=a(t_{0}),\dots$
@@ 3 A second, infinitesimally close light ray
    If we pretend to know the metric and it has nice properties, then the redshift allows us to compute the time of flight $t_{of}$ of the photon.
  • $T_{D}$ is the tracking time during which the observer tracks the source on the celestial sphere. Of course, a change in direction can only be measured with respect to another direction that supposedly does not change. Axial Bianchi I universes have indeed the luxury of such fixed directions: the $z$-axis and the $xy$-plane.
@@
$\vec{V}=\vcentcolon a_{0}\,\vec{v}\,,\quad\quad V<10\,\%.$ (9)
@@
We suppose that $T_{D}$ is small with respect to the time of flight $t_{of}$; so small that we may keep only linear terms in $T$ and $T_{D}$. Also the peculiar velocity is small enough so that we may keep only linear terms in $V$. In particular, we neglect the dilation of the observer’s proper time with respect to cosmic time $t$.
@@
Our task is to slightly change the initial direction of the second geodesic $\tilde{\vec{x}}(t)$ coded in the Noether constants $\tilde{A},\,\tilde{B}$ and $\tilde{C}$, in such a manner that $\tilde{\vec{x}}_{\tilde{0}}=\vec{x}_{0}+T_{D}\,\vec{v}$. To this end we make the uncommitted Ansatz:
$\tilde{A}=\vcentcolon A+\alpha N,\quad\tilde{B}=\vcentcolon B+\beta N,\quad\tilde{C}=\vcentcolon C+\gamma N,\ \ \text{ implying }\ \ \alpha,\,\beta,\,\gamma\sim T_{D}/t_{of}\,\vcentcolon=\,\epsilon.$ (12)
Thanks to the invariance under rescaling explained four lines after equation (6) we may choose $\tilde{N}=N$ implying
@@
(One might think that the details of equations (17) and (18) are redundant, but they will be used to compute equations (26) and (21) below.) Then we have to first order in $T/(t_{0}-t_{e})$:
$\tilde{x}_{\tilde{0}}\!\!\!$ $\sim$ $\!\!\!\int_{t_{e}+T}^{t_{0}+T_{D}}\left(\frac{A\tilde{W}}{a^{2}}\,+\,\frac{NW}{a^{2}}\,\alpha\right)$ (20)
@@
$=$ $\!\!\!x_{0}+A\,F\,T_{D}+\left(\frac{AC}{A^{2}+B^{2}}\,\gamma+\alpha\right)I-\,\frac{AC}{A^{2}+B^{2}}\,\gamma J\ +\ O(\epsilon^{2}).$ (22)
@@
$\underline{z}_{0}+1\vcentcolon=\,\frac{T_{D}}{T}\,\,=\,\frac{1}{W_{e}}\,\frac{1}{1-(A/N\,V_{x}+C/N\,V_{z})}\,+\ O(\epsilon_{s}),$ (28)
since here $T_{D}$ is an atomic period and $V<1$ may be finite. From now on we ignore $\epsilon_{s}$.
@@ 4 Arriving angles
Our first task is to compute the angle $\theta$ between the $z$-axis and the velocity $\vec{x}^{\prime}_{0}$ of the first photon. This calculation takes place in the tangent space to space-time at $(t_{0},\vec{x}_{0})$ with diagonal Lorentz-metric and evident reduction to 3-space. The negative of the spatial part of the metric (2) makes this 3-space Euclidean with the standard metric, because $a_{0}=c_{0}$. Note that the velocity vector $\vec{x}^{\prime}$ is a unit vector with respect to this 3-metric, because we have set the speed of light to one. If we denote by
@@
${\vec{x}_{e}}$.${\vec{u}_{z0}}$..${\vec{u}_{z\tilde{0}}}$.${\vec{x}(t)}$${\tilde{\vec{x}}(t)}$${{\vec{x}}^{\prime}_{0}}$${{\tilde{\vec{x}}}^{\prime}_{\tilde{0}}}$${\theta}$${\tilde{\theta}}$${\vec{v}\,T_{D}}$
Figure 1: Two infinitesimally close geodesics
Our second task is to compute the same angle $\tilde{\theta}$ for the second photon at its arrival time $t_{\tilde{0}}=t_{0}+T_{D}$ (see figure 1):
@@
$\,\frac{\tilde{C}\tilde{W}}{c}\,\sim\,\frac{CW}{c}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW}{c}\right\}^{2}\right)\gamma\right],$ (36)
@@
$\,\frac{\tilde{C}\tilde{W}_{\tilde{0}}}{c_{\tilde{0}}}\,\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW_{0}}{c_{0}}\right\}^{2}\right)\gamma\right]\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N}{C}\,\gamma\right].$ (37)
@@
Note that if $A=B=0$ the first photon moves along the $z$-axis, $\sin\theta=0$. Then the projection of the photon’s velocity onto the $xy$-plane vanishes and the angle $\varphi$ is not defined. This is the well-known coordinate singularity of spherical coordinates at the poles.
@@ 5 Linearizing the scale factors
We know that the anisotropy of the universe cannot be large and we will concentrate on axial Bianchi I universes close to an underlying maximally symmetric one with only one scale factor $a_{F}(t)$, $\cdot_{F}$ for Friedman. It is convenient to parameterize the small deviation $\eta(t)$ in the following way:
$a=\vcentcolon a_{F}\,[1-{\textstyle\frac{1}{2}}\eta],\quad c=\vcentcolon a_{F}\,[1+\eta].$ (50)
Other linearizations are viable in Bianchi I universes [13]. This is not true in the two other axial universes that admit maximal symmetry in the limit of small $\eta\,$: Bianchi V and IX. There the above linearization is mandatory [12]. (As we are presently working on axial Bianchi IX universes and their flat limit, axial Bianchi I, I choose this parameterization here.)
Indicating the linearization in $\eta$ and $V$ (neglecting cross-terms $\eta V$) also by $\sim$ we have,
$\eta_{0}=0,\quad H\sim H_{F}-{\textstyle\frac{1}{2}}\eta^{\prime},\quad H_{c}\sim H_{F}+\eta^{\prime},$ (51)
and
@@
$\alpha$ $\sim$ $\epsilon_{d}\,(3\,\cos\varphi\,\sin\theta\,\cos^{2}\theta\,\eta_{e}+[1-\cos^{2}\varphi\,\sin^{2}\theta]\,V_{x}$ (56)
@@
$\gamma$ $\sim$ $-\sin\theta\,\epsilon_{d}\left(3\,\sin\theta\,\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right),$ (57)
@@
$\chi_{e0}\vcentcolon=\int_{t_{e}}^{t_{0}}\,\frac{1}{a_{F}}\,\quad\text{and}\quad t_{of}\sim a_{F0}\,\chi_{e0}.$ (58)
Consider first the drift with respect to $\varphi$, equation (49), and linearize it in $\eta$ and $V$, but not in $T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}\sim\epsilon_{d}$:
$\cos\tilde{\varphi}\sim\frac{\cos\varphi+[3\cos\varphi\,\cos^{2}\!\theta\,\eta_{e}+(1-\cos^{2}\!\varphi\,\sin^{2}\!\theta)\!\!\left\{V_{x}/\sin\theta\right\}-\cos\varphi\,\cos\theta\,V_{z}]\,\epsilon_{d}}{\sqrt{1+2\cos\theta[3\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\left\{V_{x}/\sin\theta\right\}-V_{z}]\,\epsilon_{d}}}\,.$ (59)
@@
$\cos\tilde{\varphi}$ $\sim$ $\cos\varphi\left[1+\,\frac{\sin^{2}\varphi}{\cos\varphi}\,\frac{V_{x}}{\sin\theta}\,\epsilon_{d}\right].$ (61)
@@
$\delta\varphi$ $\sim$ $\frac{-1}{\sin\varphi}\,\,\delta\cos\varphi\sim-\,{\sin\varphi}\,\frac{V_{x}}{\sin\theta}\,\epsilon_{d}\,.$ (63)
@@
We encounter the well-known singularity of polar coordinates at the poles: If $\sin\theta=0$, then our source lies on the $z$-axis, there is no $\theta$-drift and the azimuthal angle $\varphi$ is ill-defined. Now if $\sin\theta$ is small, our perturbative parameter $V_{x}/\sin\theta$ becomes large and the linear approximation fails. That means that for sources close to the $z$-axis we cannot use the $\varphi$-drift to constrain the parameters to be fitted. We anticipate that this truncation will not reduce the statistical power of the fit significantly.
@@
$\cos\tilde{\theta}$ $\sim$ $\cos\theta\left[1-\sin^{2}\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\,\right)\epsilon_{d}\right],$ (64)
$\delta\cos\theta$ $\sim$ $-\sin^{2}\theta\,\cos\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\right)\epsilon_{d},$ (65)
$\delta\theta$ $\sim$ $\left(3\sin\theta\,\cos\theta\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right)\epsilon_{d}\,.$ (66)
@@
Note that the scale factors today $a_{F0},\,a_{0}$ and $c_{0}$ are non-essential parameters and may be chosen arbitrarily. We follow the usual convention and set them equal implying $\eta(t_{0})=0$. On the other hand $\eta^{\prime}_{0}\vcentcolon=\eta^{\prime}(t_{0})$ is an essential parameter also called ‘Hubble stretch’.
@@
In contrast to Friedman universes (axial) Bianchi I universes are not isotropic. This anisotropy manifests itself via a direction-dependent Lemaître-Hubble ‘constant’ (quantified by the Hubble stretch $\eta^{\prime}_{0}$) as well as via a direction-dependent redshift for fixed emission time (quantified by $\eta_{e}$), equation (53). The above equation (69) tells us that already for a redshift $\underline{z}_{0}=1$ the drift is hugely dominated by the second manifestation.
This is in contradiction with the result by Quercellini et al. [6][7]. They claim that – as the redshift ranges from zero to one – the absolute value of the ratio is at most 7 %. They also find the opposite sign in front of the first term.
Our independent calculation confirms the result by Marcori et al. [10] restricted to the axial case and to a comoving source.